OpenStudy (anonymous):

What is the measure of ABC A= 46° C= 58° B=? A 104° B 76° C 44° D 86°

OpenStudy (anonymous):

ABC is a triangle! And the sum of all the angles in a triangle is 180° !

OpenStudy (anonymous):

So A + B + C = 180 Now plug in A = 46, and C = 58 into it and solve for B! 46 + B + 58 = 180 Solve for B! :)

OpenStudy (anonymous):

So do I add 46 + 58 = ? ?-180

jabez177 (jabez177):

Yes, 46 + 58 = an answer then you do 180 - that answer

OpenStudy (anonymous):

You would add 46 and 58 and then subtract that amount from 180 to get your answer.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

76 should be your final answer @Pleasehelpformedal

OpenStudy (anonymous):

@Smaudri00 Giving the answer won't help them learn, now will it? Please let the asker do some of the work so they can understand the question! :) Also, welcome to OpenStudy! You might want to get to know some of the rules: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

I walked her through the question, I did not give her the answer out of nowhere... but thx for the advice anyways :) You really helped me! :)

OpenStudy (anonymous):

@MeganXOXO

OpenStudy (anonymous):

@MeganXOXO "I will encourage and guide those needing help, and not just give them an answer" I guided those in need of help, I did not JUST give her an answer. :)

OpenStudy (anonymous):

:) But from next time, keep in mind that the asker should do some work. If you tell them the steps and give them the answer. More likely then not, they will just take the answer and thank you. They won't learn anything, which will defeat the purpose :(

OpenStudy (anonymous):

.gotchya.